OpenStudy (anonymous):

i need help with this question:Emma and Rita are at a florist shop. They are planning to make bouquets of flowers for their grandmothers out of roses and carnations. The cost for a single rose is r, and the cost for a single carnation is c. Emma will use 6 carnations and 4 roses for her bouquet, and Rita will use 5 carnations and 6 roses. Write expressions that represent the total cost of Emma’s bouquet, the total cost of Rita’s bouquet, and the total cost for both bouquets combined.

OpenStudy (anonymous):

cost = (number of items)(cost of each one) for Emma 4r + 6c = emmas total cost for Rita 6r + 5c = Ritas total cost

OpenStudy (anonymous):

and for both of them combined it is emmas total cost + ritas total cost = 4r + 6c + 6r + 5c = 10r + 11c = total cost for both bouquets